OpenStudy (anonymous):

U-substitution for definite integrals. Can someone please check where I went wrong - answer supposed to be 3...

OpenStudy (anonymous):

\[\int\limits_{0}^{13}\frac{ dx }{ \sqrt[3]{(1+2x)^2} }\] \[u=1+2x\] \[du=2dx\] changing limits of integration I get \[\frac{ 1 }{ 2 } \int\limits_{1}^{27}u^\frac{ -2 }{ 3 }du\] \[=\frac{ 1 }{ 2 }(-3u^\frac{ -1 }{ 3 } )\] \[(-\frac{ 3 }{ 2\sqrt[3]{27} }+\frac{ 3 }{ 2\sqrt[3]{1} })=-\frac{ 1 }{ 2 }+\frac{ 3 }{ 2 }=1\]

OpenStudy (cwrw238):

should n't your penultimate line be 1/2(3 u ^ (1/3) ?

OpenStudy (science0229):

When you were integrating u^-2/3, it should be 3u^1/3, not 3u^-1/3

OpenStudy (cwrw238):

yep that's it

OpenStudy (anonymous):

I see it, thanks!

OpenStudy (science0229):

Yep.